OpenStudy (anonymous):

Find (3y-1)^2 please help because i am not that good at math and i dont have much time to complete this

OpenStudy (anonymous):

please show your work if possible thank you and God bless!

OpenStudy (solomonzelman):

You can either use \(\large \color{blue}{ \bf (a-b)^2=a^2-2ab+b^2 }\) or pascal's Triangle \(\Large\color{red}{ \bf 1 }\) \(\Large\color{red}{ \bf 1~~~~~~~~1 }\) \(\Large\color{blue}{ \bf 1~~~~~~~2~~~~~~~1 }\) \(\Large\color{red}{ \bf 1~~~~~~3~~~~~~~~3~~~~~~~1 }\) \(\Large\color{red}{ \bf 1~~~~~~4~~~~~~~6~~~~~~~4~~~~~~~1 }\) \(\Large\color{red}{ \bf 1~~~~~5~~~~~10~~~~~~10~~~~~~5~~~~~~1 }\)

OpenStudy (anonymous):

i havent learned pascals triangle yet, so i am guessing it would be 3^2-2(3)(1)+1^2 is that correct?

OpenStudy (solomonzelman):

well (3y)^2-2(3y)1+1^2 don't forger the "y". Now , expand it...

OpenStudy (anonymous):

hello?

OpenStudy (ranga):

It would be: \( (3y)^2 - 2(3y)(1) + (1)^2\)

OpenStudy (anonymous):

so thats the final answer or is it possible to simplify it

OpenStudy (ranga):

Simplify it. \((3y)^2 - 2(3y)(1) + (1)^2 = 9y^2 - 6y + 1\)

OpenStudy (anonymous):

oh ok thank you

OpenStudy (ranga):

You are welcome.